# FutureSightDrive: Thinking Visually with Spatio-Temporal CoT for Autonomous Driving

### v2

@@ 3.4 Training strategy
Our objective is to preserve understanding capabilities of MLLMs through VQA tasks and activate their visual generation capabilities to predict future frames. VQA task data originates from OmniDrive-nuScenes [64]. We incorporate a large volume of unlabeled image data from nuScenes [1] for future frame prediction. To implement progressive easy-to-hard CoT, we integrate nuScenes annotated data to teach the model predicting image-formatted future lane dividers and 3D detection. Finally, we add future frame prediction with CoT datas containing intermediate reasoning steps. All the above understanding and generation tasks are trained together.
@@
We focus on autonomous driving scene understanding and trajectory planning. Following OmniDrive [64], scene understanding utilizes DriveLM’s GVQA [54] dataset. For trajectory planning, we follow VAD [29, 21] using nuScenes, where our spatio-temporal CoT integrates the holistic future scene, explicit lane dividers, and 3D detection results into a single future frame as intermediate reasoning steps. We train these tasks simultaneously using a single model, enabling task-specific predictions during inference through different task prompts.
@@ 4.1 Experimental settings
Following the previous methods [29, 13, 4], we evaluate trajectory planning and future frames generation on the nuScenes [1]. The nuScenes contains 1,000 scenes of approximately 20 seconds each captured by a 32-beam LiDAR and six cameras providing 360-degree field of view. Specifically, The dataset provides 28,130 (train), 6,019 (val), and 193,082 (unannotated) samples. Additionally, we conducted experiments on NAVSIM [10], a realistic scenario dataset designed for real-world planning. This dataset aims to highlight challenging driving scenarios involving dynamic changes in driving intent, while deliberately excluding simple situations such as static scenes or constant-speed driving.
Following the previous methods [7, 64], we evaluate scene understanding on DriveLM [54]. This dataset features keyframe descriptions paired with QA annotations covering full-stack autonomous driving (perception, prediction, planning), offering comprehensive language support for development.
@@
We evaluate trajectory planning using L2 displacement error and collision rate following previous methods [21, 29, 19]. Notably, UniAD [21] computes L2 metrics and collision rate at each timestep, whereas ST-P3 [19] and VAD [29] considers the average of all previous time-steps. For a fair comparison, we adopted these two different calculation methods. Following existing methods [65, 77, 71], we report Fréchet Inception Distance (FID) [17] to measure the future frames generation quality. DriveLM GVQA [54] metrics include language metrics like BLEU, ROUGE_L, and CIDEr for text generation, the ChatGPT Score for open-ended Q&A and accuracy for multiple-choice questions. For NAVSIM [10], we adopt the official metrics for evaluation, especially PDMS.
@@
We initialize our model with Qwen2-VL-2B [63] and pre-train it for 32 epochs to enable visual generation while preserving semantic understanding. During fine-tuning (12 epochs on 8 NVIDIA RTX A6000), we use $1\times 10^{-4}$ learning rate and batch size of 16. We expand the visual codebook of MoVQGAN [92] to the vocabulary of the large language model and use its detokenizer to convert the visual tokens predicted by the large language model to the pixel space.
Table 1: End-to-end trajectory planning experiments on nuScenes [1]. We evaluated the L2 and collision metrics based on the distinct computational methodologies of ST-P3 [19] and UniAD [21], respectively. * indicates that the ego status is additionally used. VAD [29] and UniAD [21] results are derived from BEV-Planner [34], while the remaining results are sourced from their respective papers.
@@
ST-P3* [ECCV22] [19] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - - -
VAD [ICCV23] [29] 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09 - - - - - - - - -
VAD* [ICCV23] [29] 0.17 0.34 0.60 0.37 0.04 0.27 0.67 0.33 - - - - - - - - -
UniAD [CVPR23] [21] - - - - - - - - 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77 -
UniAD* [CVPR23] [21] - - - - - - - - 0.20 0.42 0.75 0.46 0.02 0.25 0.84 0.37 -
BEV-Planner [CVPR24] [34] 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59 - - - - - - - - -
BEV-Planner* [CVPR24] [34] 0.16 0.32 0.57 0.35 0.00 0.29 0.73 0.34 - - - - - - - - -
PreWorld [ICLR25] [32] - - - - - - - - 0.49 1.22 2.32 1.34 0.19 0.57 2.65 1.14 -
@@
ELM [ECCV24] [98] - - - - - - - - 0.34 1.23 2.57 1.38 0.12 0.50 2.36 0.99 BLIP2-2.7B
FeD* [CVPR24] [86] - - - - - - - - 0.27 0.53 0.94 0.58 0.00 0.04 0.52 0.19 LLaVA-7B
OccWorld [ECCV24] [94] 0.39 0.73 1.18 0.77 0.11 0.19 0.67 0.32 0.52 1.27 2.41 1.40 0.12 0.40 2.08 0.87 GPT3-like
Doe-1 [arxiv24] [95] 0.37 0.67 1.07 0.70 0.02 0.14 0.47 0.21 0.50 1.18 2.11 1.26 0.04 0.37 1.19 0.53 Lumina-mGPT-7B
RDA-Driver* [ECCV24] [26] 0.17 0.37 0.69 0.40 0.01 0.05 0.26 0.10 0.23 0.73 1.54 0.80 0.00 0.13 0.83 0.32 LLaVA-7B
EMMA* [arxiv24] [27] 0.14 0.29 0.54 0.32 - - - - - - - - - - - - Gemini 1-1.8B
OmniDrive [CVPR25] [64] 0.40 0.80 1.32 0.84 0.04 0.46 2.32 0.94 - - - - - - - - LLaVA-7B
OmniDrive* [CVPR25] [64] 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30 - - - - - - - - LLaVA-7B
@@
VADv2 [arXiv24] [3] 97.2 89.1 91.6 100 76.0 80.9
UniAD [CVPR23] [21] 97.8 91.9 92.9 100 78.8 83.4
DiffusionDrive-Cam [CVPR25] [36] 97.8 92.2 92.6 99.9 78.9 83.6
LTF [TPAMI23] [6] 97.4 92.8 92.4 100 79.0 83.8
PARA-Drive [CVPR24] [69] 97.9 92.4 93.0 99.8 79.3 84.0
LAW [ICLR25] [33] 96.4 95.4 88.7 99.9 81.7 84.6
@@ 4.2 Main results
We present trajectory planning performance on nuScenes following previous methods [29, 21] in Table 1. When using ego status, FSDrive surpasses previous SOTA methods using ego status in ST-P3 and UniAD metrics. However, following BEV-Planner [34] findings about ego-status’s performance boost, we prioritize non-ego-status evaluations. Compared to non-autoregressive (e.g., UniAD) and autoregressive methods (e.g., OmniDrive), FSDrive demonstrates superior effectiveness. Notably, FSDrive outperforms Doe-1 [95] which also enables vision generation (L2: 0.53 vs. 0.70 and 0.96 vs. 1.26; collision: 0.19 vs. 0.21 and 0.40 vs. 0.53), indicating limitations in their VQ-VAE-based discrete visual features for understanding. For a fair comparison, we also used LLaVA like methods [64, 26, 86, 75]. Under the corresponding settings, FSDrive still has excellent competitiveness, indicating that FSDrive can be widely applied to any existing MLLM.
Table 3: Future frames generation results on the nuScenes [1] dataset.
@@
 [CVPR21 [30]]  [ECCV24 [65]]  [CVPR24 [66]]  [CVPR24 [77]]  [CVPR25 [16]]  [arxiv24 [95]]
@@
Table 2 shows the evaluation results for NAVSIM [10]. All approaches rely exclusively on camera input, with no lidar data being used. Achieving a PDMS score of 85.1, FSDrive outperforms prior camera-only methods like LAW [33] and DiffusionDrive-Cam [36], thus showcasing its efficacy in the pseudo closed-loop setting.
@@
FSDrive’s scene understanding was evaluated on DriveLM in Table 4, achieving 0.57 and outperforming recent methods like Cube-LLM [7] and OmniDrive [64]. This highlights the effectiveness of FSDrive pre-training paradigm for generation and understanding.
Table 4: Results on DriveLM [54] GVQA benchmark.
@@
DriveLM baseline [54] 0.00 0.65 0.05 0.08 0.10 0.28 0.32
Cube-LLM [7] 0.39 0.89 0.16 0.20 0.31 0.39 0.50
TrackingMeetsLMM [28] 0.60 0.58 0.72 0.72 0.04 0.36 0.52
SimpleLLM4AD [93] 0.66 0.57 0.76 0.73 0.15 0.35 0.53
OmniDrive [64] 0.70 0.65 0.52 0.73 0.13 0.37 0.56
@@ 4.3 Ablation study
In this section, unless otherwise specified, we evaluate the computing metrics of UniAD [21] based on the Qwen2-VL-2B model [63] and do not use the ego status.
@@
None 29.4
$\sim$100k 16.2
$\sim$200k 12.7
(b) Ablations of progressive method.
[…]
Method
$\times$ 12.7
$\checkmark$ 10.1
@@
We conduct ablation studies on future frames generation in Table 7(b). Table 7(a) shows that larger pre-training datasets improve MLLM’s visual generation capability. Despite being much smaller (200K vs. 100M in previous work [70]), our data achieves more robust visual generation. Scaling datasets may further enhance performance. Table 7(b) confirms our progressive method improves autoregressive image generation.

### v3

@@ 3.4 Training strategy
Our objective is to preserve understanding capabilities of MLLMs through VQA tasks and activate their visual generation capabilities to predict future frames. VQA task data originates from OmniDrive-nuScenes [Wang2024OmniDriveAH]. We incorporate a large volume of unlabeled image data from nuScenes [Caesar2019nuScenesAM] for future frame prediction. To implement progressive easy-to-hard CoT, we integrate nuScenes annotated data to teach the model predicting image-formatted future lane dividers and 3D detection. Finally, we add future frame prediction with CoT datas containing intermediate reasoning steps. All the above understanding and generation tasks are trained together.
@@
We focus on autonomous driving scene understanding and trajectory planning. Following OmniDrive [Wang2024OmniDriveAH], scene understanding utilizes DriveLM’s GVQA [sima2023drivelm] dataset. For trajectory planning, we follow VAD [jiang2023vad, hu2023_uniad] using nuScenes, where our spatio-temporal CoT integrates the holistic future scene, explicit lane dividers, and 3D detection results into a single future frame as intermediate reasoning steps. We train these tasks simultaneously using a single model, enabling task-specific predictions during inference through different task prompts.
@@ 4.1 Experimental settings
Following the previous methods [jiang2023vad, gao2024vista, chen2025technical], we evaluate trajectory planning and future frames generation on the nuScenes [Caesar2019nuScenesAM]. The nuScenes contains 1,000 scenes of approximately 20 seconds each captured by a 32-beam LiDAR and six cameras providing 360-degree field of view. Specifically, The dataset provides 28,130 (train), 6,019 (val), and 193,082 (unannotated) samples. Additionally, we conducted experiments on NAVSIM [navsim], a realistic scenario dataset designed for real-world planning. This dataset aims to highlight challenging driving scenarios involving dynamic changes in driving intent, while deliberately excluding simple situations such as static scenes or constant-speed driving.
Following the previous methods [cube-llm, Wang2024OmniDriveAH], we evaluate scene understanding on DriveLM [sima2023drivelm]. This dataset features keyframe descriptions paired with QA annotations covering full-stack autonomous driving (perception, prediction, planning), offering comprehensive language support for development.
@@
We evaluate trajectory planning using L2 displacement error and collision rate following previous methods [hu2023_uniad, jiang2023vad, hu2022stp3]. Notably, UniAD [hu2023_uniad] computes L2 metrics and collision rate at each timestep, whereas ST-P3 [hu2022stp3] and VAD [jiang2023vad] considers the average of all previous time-steps. For a fair comparison, we adopted these two different calculation methods. Following existing methods [wang2023drivedreamer, yang2024genad, 3DGAN], we report Fréchet Inception Distance (FID) [fid] to measure the future frames generation quality. DriveLM GVQA [sima2023drivelm] metrics include language metrics like BLEU, ROUGE_L, and CIDEr for text generation, the ChatGPT Score for open-ended Q&A and accuracy for multiple-choice questions. For NAVSIM [navsim], we adopt the official metrics for evaluation, especially PDMS.
@@
We initialize our model with Qwen2-VL-2B [Qwen2-VL] and pre-train it for 32 epochs to enable visual generation while preserving semantic understanding. During fine-tuning (12 epochs on 8 NVIDIA RTX A6000), we use $1\times 10^{-4}$ learning rate and batch size of 16. We expand the visual codebook of MoVQGAN [Zheng2022MoVQgan] to the vocabulary of the large language model and use its detokenizer to convert the visual tokens predicted by the large language model to the pixel space.
Table 1: End-to-end trajectory planning experiments on nuScenes [Caesar2019nuScenesAM]. We evaluated the L2 and collision metrics based on the distinct computational methodologies of ST-P3 [hu2022stp3] and UniAD [hu2023_uniad], respectively. * indicates that the ego status is additionally used. VAD [jiang2023vad] and UniAD [hu2023_uniad] results are derived from BEV-Planner [bev-planner], while the remaining results are sourced from their respective papers.
@@
ST-P3* [ECCV22] [hu2022stp3] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - - -
VAD [ICCV23] [jiang2023vad] 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09 - - - - - - - - -
VAD* [ICCV23] [jiang2023vad] 0.17 0.34 0.60 0.37 0.04 0.27 0.67 0.33 - - - - - - - - -
UniAD [CVPR23] [hu2023_uniad] - - - - - - - - 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77 -
UniAD* [CVPR23] [hu2023_uniad] - - - - - - - - 0.20 0.42 0.75 0.46 0.02 0.25 0.84 0.37 -
BEV-Planner [CVPR24] [bev-planner] 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59 - - - - - - - - -
BEV-Planner* [CVPR24] [bev-planner] 0.16 0.32 0.57 0.35 0.00 0.29 0.73 0.34 - - - - - - - - -
PreWorld [ICLR25] [li2025semi] - - - - - - - - 0.49 1.22 2.32 1.34 0.19 0.57 2.65 1.14 -
@@
ELM [ECCV24] [zhou2024embodied] - - - - - - - - 0.34 1.23 2.57 1.38 0.12 0.50 2.36 0.99 BLIP2-2.7B
FeD* [CVPR24] [zhang2023coaching] - - - - - - - - 0.27 0.53 0.94 0.58 0.00 0.04 0.52 0.19 LLaVA-7B
OccWorld [ECCV24] [zheng2023occworld] 0.39 0.73 1.18 0.77 0.11 0.19 0.67 0.32 0.52 1.27 2.41 1.40 0.12 0.40 2.08 0.87 GPT3-like
Doe-1 [arxiv24] [doe] 0.37 0.67 1.07 0.70 0.02 0.14 0.47 0.21 0.50 1.18 2.11 1.26 0.04 0.37 1.19 0.53 Lumina-mGPT-7B
RDA-Driver* [ECCV24] [RDA-Driver] 0.17 0.37 0.69 0.40 0.01 0.05 0.26 0.10 0.23 0.73 1.54 0.80 0.00 0.13 0.83 0.32 LLaVA-7B
EMMA* [arxiv24] [Hwang2024EMMAEM] 0.14 0.29 0.54 0.32 - - - - - - - - - - - - Gemini 1-1.8B
OmniDrive [CVPR25] [Wang2024OmniDriveAH] 0.40 0.80 1.32 0.84 0.04 0.46 2.32 0.94 - - - - - - - - LLaVA-7B
OmniDrive* [CVPR25] [Wang2024OmniDriveAH] 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30 - - - - - - - - LLaVA-7B
@@
VADv2 [arXiv24] [chen2024vadv2] 97.2 89.1 91.6 100 76.0 80.9
UniAD [CVPR23] [hu2023_uniad] 97.8 91.9 92.9 100 78.8 83.4
DiffusionDrive-Cam [CVPR25] [diffusiondrive] 97.8 92.2 92.6 99.9 78.9 83.6
LTF [TPAMI23] [ltf] 97.4 92.8 92.4 100 79.0 83.8
PARA-Drive [CVPR24] [Weng2024paradrive] 97.9 92.4 93.0 99.8 79.3 84.0
LAW [ICLR25] [law] 96.4 95.4 88.7 99.9 81.7 84.6
@@ 4.2 Main results
We present trajectory planning performance on nuScenes following previous methods [jiang2023vad, hu2023_uniad] in Table 1. When using ego status, FSDrive surpasses previous SOTA methods using ego status in ST-P3 and UniAD metrics. However, following BEV-Planner [bev-planner] findings about ego-status’s performance boost, we prioritize non-ego-status evaluations. Compared to non-autoregressive (e.g., UniAD) and autoregressive methods (e.g., OmniDrive), FSDrive demonstrates superior effectiveness. Notably, FSDrive outperforms Doe-1 [doe] which also enables vision generation (L2: 0.53 vs. 0.70 and 0.96 vs. 1.26; collision: 0.19 vs. 0.21 and 0.40 vs. 0.53), indicating limitations in their VQ-VAE-based discrete visual features for understanding. For a fair comparison, we also used LLaVA like methods [Wang2024OmniDriveAH, RDA-Driver, zhang2023coaching, xie2025seqgrowgraph]. Under the corresponding settings, FSDrive still has excellent competitiveness, indicating that FSDrive can be widely applied to any existing MLLM.
Table 3: Future frames generation results on the nuScenes [Caesar2019nuScenesAM] dataset.
@@
 [CVPR21 [kim2021drivegan]]  [ECCV24 [wang2023drivedreamer]]  [CVPR24 [wang2023driving]]  [CVPR24 [yang2024genad]]  [CVPR25 [Hassan2024GEMAG]]  [arxiv24 [doe]]
@@
Table 2 shows the evaluation results for NAVSIM [navsim]. All approaches rely exclusively on camera input, with no lidar data being used. Achieving a PDMS score of 85.1, FSDrive outperforms prior camera-only methods like LAW [law] and DiffusionDrive-Cam [diffusiondrive], thus showcasing its efficacy in the pseudo closed-loop setting.
@@
FSDrive’s scene understanding was evaluated on DriveLM in Table 4, achieving 0.57 and outperforming recent methods like Cube-LLM [cube-llm] and OmniDrive [Wang2024OmniDriveAH]. This highlights the effectiveness of FSDrive pre-training paradigm for generation and understanding.
Table 4: Results on DriveLM [sima2023drivelm] GVQA benchmark.
@@
DriveLM baseline [sima2023drivelm] 0.00 0.65 0.05 0.08 0.10 0.28 0.32
Cube-LLM [cube-llm] 0.39 0.89 0.16 0.20 0.31 0.39 0.50
TrackingMeetsLMM [ishaq2025trackingmeets] 0.60 0.58 0.72 0.72 0.04 0.36 0.52
SimpleLLM4AD [Zheng2024SimpleLLM4ADAE] 0.66 0.57 0.76 0.73 0.15 0.35 0.53
OmniDrive [Wang2024OmniDriveAH] 0.70 0.65 0.52 0.73 0.13 0.37 0.56
@@ 4.3 Ablation study
In this section, unless otherwise specified, we evaluate the computing metrics of UniAD [hu2023_uniad] based on the Qwen2-VL-2B model [Qwen2-VL] and do not use the ego status.
@@
None $\times$ 29.4
$\sim$100k $\times$ 16.2
$\sim$200k $\times$ 12.7
$\sim$200k $\checkmark$ 10.1
@@
We conduct ablation studies on future frames generation in Table 7. The upper part of Table 7 shows that larger pre-training datasets improve MLLM’s visual generation capability. Despite being much smaller (200K vs. 100M in previous work [wu2024janus]), our data achieves more robust visual generation. Scaling datasets may further enhance performance. The lower part of Table 7 confirms our progressive method improves autoregressive image generation.